OpenStudy (deleonbrrj):

How fast must a ball be thrown vertically from the bottom up to reach a height of 50m? How long it will remain in the air?

OpenStudy (anonymous):

remeber the kinematics, you will need two of the equations for this one: first: vf^2=vi^2+a*d then once you find vi, plug it into the second: vf=vi+a*t when throwing a ball up, when it reaches its max height, the final velocity is 0. and the distance we are looking for is 50, and acceleration due to gravity is -9.8m/s^2 so: 0=vi^2+(-9.8m/s^2)*50m so -vi^2=-490 m^2/s^2 vi=sqrt(490m^2/s^2) vi=22.14m/s then plug it in: vf=vi+a*t 0=22.14m/s+(-9.8m/s^2)*t -22.14=-9.8*t -22.14m/s/-9.8m/s^2=t t=2.26s ** don't hate me if it's wrong :(( **

OpenStudy (anonymous):

um...i think ur formula is wrong im pretty sure it is vf^2=vo^2+2ad

OpenStudy (anonymous):

otherwise i think ur thought process is correct

OpenStudy (anonymous):

i got initial speed 31.32m/s and time 6.38s

OpenStudy (deleonbrrj):

Perfect Cherio12, i got this result now tbem

OpenStudy (anonymous):

no problem